OpenStudy (anonymous):

θ is any real number. is it possible for sin θ = 5/4? why or why not?

OpenStudy (agent0smith):

sinθ is always... \[-1 \le \sin \theta \le 1\]

OpenStudy (anonymous):

thanks :) so is it possible in that case?

jimthompson5910 (jim_thompson5910):

basically agent0smith is saying sin θ is between -1 and 1 for any value of θ

jimthompson5910 (jim_thompson5910):

5/4 = 1.25 so what do you think? yes? no?

OpenStudy (anonymous):

ohh ok i see

OpenStudy (anonymous):

so no it is not possible

jimthompson5910 (jim_thompson5910):

ok great yes, it's not possible

OpenStudy (anonymous):

can you help me with one more like this? i'd really appreciate it.. i'm kind of confused. i'll post now :)